OpenStudy (anonymous):

What is the solution to the equation 9^(x + 1) = 27 ?

OpenStudy (anonymous):

x = 2.5 x = 0.5 x = –0.5 x = –3.5

OpenStudy (zale101):

takes logs on both sides \[\log 9^{(x+1)}=\log 27\] \[(x+1)\log9=\log 27\] move the power to the front \[x+1=\frac{ \log27 }{ \log 9 }\] divide the like terms, logs \[x+1=1.5\] subtract 1 to both sides x=0.5

OpenStudy (anonymous):

Thank You! I was really lost.

OpenStudy (zale101):

no problem :)

hartnn (hartnn):

\(x+1=\dfrac{ \log27 }{ \log 9 } = \dfrac{\log 3^3}{\log 3^2} = \dfrac{3 \log 3}{2\log 3} = \dfrac{3}{2}=1.5\)

OpenStudy (zale101):

a good way of extending it hartnn :)